THE INFRARED SPECTRUM OF CH_{3}O IN THE CH STRETCHING REGION

JIAXIANG HAN^{a} and R. F. CURL, Rice University, Department of Chemistry and Rice Quantum Institute, Houston, TX 77005; \underline{SHUIMING HU}, University of Science and Technology of China, Hefei, PRC; J. M. BROWN, Physical and Theoretical Chemistry Laboratory, Oxford University, Oxford OX1 3QZ, United Kingdom.

Since the last report,^{b} we have made progress in obtaining and analyzing the jet-cooled high resolution spectrum of methoxy, CH_{3}O, in the CH stretching region. Specifically, a second perpendicular band similar to the first in exhibiting Δp=+1 has been obtained and fitted. In addition two additional isolated subbands, one with perpendicular selection rules and the other with parallel selection rules have been assigned and fitted. We are attempting to achieve a global understanding of these results.

aPresent address: Department of Chemistry, University of California, Berkeley, CA 94720.

bJia-xiang Han, Yu. G. Utkin, Hong-bing Chen, L. A. Burns, and R. F. Curl, J. Chem. Phys. 117, 6538-6545 (2002).